Mark had an equal number of blue and red marbles. He gave away 5/9 of his
blue marbles and sold 49 red marbles. In the end, he was left with 1/3 of the
marbles. How many marbles did he have at first?

Answer by VFBundy(438)

Number of red marbles (at start) = n
Number of blue marbles (at start) = n

Number of marbles remaining after giveaways and sales:

4/9(n) + (n - 49) = 1/3(2n)

4/9(n) + (n - 49) = 2/3(n)

(n - 49) = 2/3(n) - 4/9(n)

-49 = 2/3(n) - 4/9(n) - n

-49 = 6/9(n) - 4/9(n) - 9/9(n)

-49 = -7/9(n)

-49(-9/7) = n

n = 63

Number of red marbles (at start) = n = 63
Number of blue marbles (at start) = n = 63

Mark had a total of 126 marbles at the start.